OpenStudy (anonymous):

Evaluate the expression when x = 7 and y = 2 x – y – 3 A. 2 B. 5 C. 6 D. 8

OpenStudy (***[isuru]***):

just plug them in to the expression 7 - 2 - 3 = ?

OpenStudy (anonymous):

2

OpenStudy (***[isuru]***):

yep!

OpenStudy (anonymous):

Evaluate the expression when x = 8 and y = 1. A. 4 B. 8 C. D.

OpenStudy (lieutenantgeneral):

A. 2 cause 7 - 3 = 4 - 2 = 2

OpenStudy (***[isuru]***):

where is the expression ? is it same as the last one ?

OpenStudy (***[isuru]***):

x - y - 3 ?

OpenStudy (anonymous):

1/4and 1/8

OpenStudy (anonymous):

so

OpenStudy (***[isuru]***):

x - y - 3 = 8 - 1 - 3 = ?

OpenStudy (anonymous):

4

OpenStudy (***[isuru]***):

yep

OpenStudy (anonymous):

So now when I sumbiit this will I have it right?

OpenStudy (***[isuru]***):

yeah! y not :)